SS Company provides you with the following condensed balance sheet information: Assets (5) Total stockholders equity Current assets Equity investments Equipment (net) Intangibles 43,900 64,BOO 238,100 5B,300 $405,100 (c) akmuas declares and issues a 30% stock dividend when the market price of the stock is S 14 per share. Total assets (1) Total assets Liabilities and Stockholders Equity Current and long-term liabilities Stockholders equity $102,000 (2) Common stock Common stock ($5 par》 Paid-in capital in excess of par Retained earnings $19,900 112,800 170,400 (3) Paid-in capital in excess of par 303,100 $405,100 (4) Retained earnings Total liabilities and stockholders equity (5) Total stockholders equity For each transaction below, indicate the dollar impact (ift any) on the following tive items: (1) total assets, (2) cammon stock, (3) paid-in capital in excess of par, (4) retained earnings, and (S) stockholders equity. (Each situation is independent.) (d) Susoan declares and distributes a property dividend.Sksoaa gives one share of its equity investment (ABC stock) for every two shares of Sucs0ag Company stock held. SkeS0R owns 10,800 shares of ABC. ABC is selling for $9 per share on the date the property dividend is dedared. (a) Skhsoca declares and pays a $0.60 per share cash dividend. (1) Total assets (1) Total assets (2) Common stock (2) Common stock (3) Paid-in capital in excess of par (3) Paid-in capital in excess of par (4) Retained earnings (4) Retained earnings (5) Total stockholders equity (5) Total stockholders equity (b) SuoAA declares and issues a 10% stock dividend when the market price of the stock is *15 per share. (e) Skrsooa, declares a 3-far-1 stock split and issues new shares (1) Total assets (1) Total assets (2) Common stock (2) Common stock (3) Paid-in capital in excess of par (3) Paid-in capital in excess of par (4) Retained eanings (4) Retained earnings (5) Total stockholders equity
Answer #1

a) Number of Shares = 19,900/5 = 3,980 Shares

Dividend = 3,980 x 0.60 = $ 2,388

  • Total Assets= Decrease of $ 2,388
  • Common Stock = No changes
  • Paid in Capital in Excess of Par = No Changes
  • Retained Earnings = Decrease of $ 2,388
  • Total Stockholders' Equity = Decrease of $ 2,388

b) Dividend = 3,980 x10% x 15 = 5,970

in which Common Stock = 398 x 5 =1,990

Pain in Capital in excess of Par = 398 x 10 = 3,980

  • Total Assets = No changes
  • Common Stock = Increase by 1,990
  • Pain Capital in Excess of Par = Increase by 3,980
  • Retained Earnings = Decrease by 5,970
  • Total Stockholders Equity= No changes.

c) Dividend = 3,980 x 30% x 14 = 16,716

in which Common Stock = 1,194 x 5 =5,970

Pain in Capital in excess of Par = 1,194 x 9 = 10,746

  • Total Assets = No changes
  • Common Stock = Increase by 5,970
  • Pain Capital in Excess of Par = Increase by 10,746
  • Retained Earnings = Decrease by 16,716
  • Total Stockholders Equity= No changes.

d)Amount of Property dividend =3,980/2 * 9 =17,910

Value of ABC Shares = 3,980/2 x 6 = 11,940

Unrealized gain = 17,910 -11,940 = 5,970

  • Total Assets = Decrease by 11,940
  • Common Stock = No changes.
  • Pain Capital in Excess of Par = No changes.
  • Retained Earnings = Decrease by 11,940
  • Total Stockholders Equity= No changes.

e)

  • Total Assets = No changes.
  • Common Stock = No changes.
  • Pain Capital in Excess of Par = No changes.
  • Retained Earnings = No changes.
  • Total Stockholders Equity= No changes.